OpenStudy (anonymous):

A roofer calculates his bid price using the formula P = 1.85s + 4.2f, where s is the area of the roof in square feet and f is the length of the fascia in feet. Find the area of the roof with 190 feet of fascia and a price of $4,148.

OpenStudy (gw2011):

P = 1.85s + 4.2f 1.85s = P - 4.2f s = (P - 4.2f)/1.85 s = (4148 - 4.2(190))/1.85 s = (4148 - 798)/1.85 s = 1810.8 square feet